OpenStudy (anonymous):

If the point (15/13,12/13) corresponds to an angle theta to the unit circle what is cot theta

OpenStudy (anonymous):

\[\frac{ 12 }{ 5}\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 12}\]

OpenStudy (anonymous):

\[\frac{ 13 }{ 5}\]

OpenStudy (anonymous):

\[\frac{ 13 }{ 12 }\]

OpenStudy (anonymous):

\[\frac{ 5 }{ 13 }\]

OpenStudy (freckles):

think you mean (5/13,12/13) anyways yeah this point is on the unit circle since (5/13)^2+(12/13)^2=1^2 holds so on the unit circle the radius,r, is 1 recall that cot(t)=cos(t)/sin(t) where cos(t)=x and sin(t)=y so cot(t)=x/y